OpenStudy (anonymous):

@mominanisar If m/n = .75, then what is 3m + 2n?

OpenStudy (zehanz):

3m+2n could still be anything. You could try this, though:\[\frac{ m }{ n }=\frac{ 3 }{ 4 } \Leftrightarrow m=\frac{ 3 }{ 4 }n\]Substitute this in place of m:\[3 \cdot \frac{ 3 }{ 4 }n+2n\]This expression has only n as variable, totalling\[\frac{ 17 }{ 4 }n=4\frac{ 1 }{ 4 }n\]